OpenStudy (love_333):

Medal + fan! If the scale factor of two similar solids is 5:13, what is the ratio of their corresponding areas and volumes? A 125:2,197 and 25:169 B 25:169 and 125:2,197 C 10:26 and 15:39 *My Guess? D 5:169 and 5:2,197

OpenStudy (love_333):

@Erak ?

OpenStudy (love_333):

C?

OpenStudy (erak):

I'm pretty sure it's b, hold up

OpenStudy (erak):

Yeah I believe it's B, I just used a two cubes, 5x5x5 and 13x13x13 and tested their SA/areas

OpenStudy (love_333):

Okay, Well I wasn't very sure of my answer. You def. seem to be better at this than me, so i'll take your word for it lol

OpenStudy (erak):

I mean SA for 5cube = 6*5^2 = 150 SA for 13cube = 6*13^2 = 1014 Ratio is 150/1014 = 0.1479 Given ratio is 25/169 = 0.1479 Same with the volume